OpenStudy (anonymous):

When raising a quoitent to a power, what is raised to the power? numerator denominator numerator and denominator Pretty sure it's the third option.

OpenStudy (anonymous):

Try it out. 2/3 to the power of 2

OpenStudy (whpalmer4):

I'm pretty sure you're right :-) One of the properties of exponents is \[(\frac{a}{b})^n = \frac{a^n}{b^n}\] (so long as a and b aren't 0)

jimthompson5910 (jim_thompson5910):

You are correct, example: (5/4)^2 = (5^2)/(4^2) (5/4)^2 = 25/16

OpenStudy (anonymous):

I see. Thanks. :)